Prove that Newton's second law is dimensionally correct
1
Expert's answer
2020-06-15T10:27:12-0400

Answer: Newton's second law is formulated as follows: the acceleration that a body acquires as a result of a force is directly proportional to that force and inversely proportional to the mass of the body.

"\overrightarrow{a}=\dfrac{\overrightarrow{F}}{m};"

1N is equal to the force which, acting on a body of mass 1kg, gives it an acceleration of 1m / s2.


"[a]=\dfrac{1N}{1kg}=\dfrac{1kg\sdot1m/s^2}{1kg}=1m/s^2."